OpenStudy (anonymous):

If p% of q is k, then q% of p is.... A) k/100 B) pq/100 C) pk/100 D) qk/100 E) k

OpenStudy (lgbasallote):

p% of q is k so \[\frac{p}{100} \times q = k\] \[\frac{pq}{100} = k\] what is q% of p q% of p would be \[\frac{q}{100} \times p \implies \frac{pq}{100}\] so what do you think is the answer?

OpenStudy (lgbasallote):

tag me if you have additional questions @Genuine

OpenStudy (anonymous):

Is it k or pq/100? Because I just got k as my answer

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

yes it's k

OpenStudy (anonymous):

because